# Scalable Multi-Robot Motion Planning Using Guidance-Informed Hypergraphs

### v2

@@ IV-C Kinodynamic Planning
Incorporating kinodynamic constraints into Workspace Guided-DaSH is straightforward due to the modularity of the DaSH framework. As kinodynamic constraints primarily influence the motion of the robots, it is imperative that the motion hypergraph construction process accounts for these constraints. To this end, we adapt the Composite-RRT and CDR-RRT algorithms, employed in the construction of the motion hypergraph in Algorithm 3, to incorporate kinodynamic constraints. This is achieved by substituting their underlying RRT with our kinodynamic RRT, following a similar approach to [24].
We model the dynamics of each robot as follows:
$\dot{x_{i}}=f_{i}(x_{i},u_{i})$ (1)
where $x_{i}\in\mathcal{X}_{i}$ represents the state of robot $i$, and $u_{i}\in\mathcal{U}_{i}$ denotes the control input of robot $i$ with $\mathcal{X}_{i}$ and $\mathcal{U}_{i}$ being the respective state space and control space of each agent.
Drawing on the work presented in [24], we outline our kinodynamic RRT in 5, with the kinodynamic steering function specified in Algorithm 6.
IV-C1 Kinodynamic RRT
Algorithm 6 begins by sampling a control input for each robot uniformly at random from their respective control space $\mathcal{U}_{i}$. The control inputs are sampled either uniformly at random from the control space or by selecting the control that brings the robot closest to the target state $x_{rand}$. To find the closest control, multiple controls are sampled uniformly at random from the robots’ control space and integrated forward in time for a set duration. The control which drives the robot state closest to $x_{rand}$ is used to find $x_{new}$. Sampling the control in this manner enables effective utilization of the guidance provided by CDR-RRT, as it ensures that the motion is directed as closely as possible toward the direction of the guiding configuration.
Next, a duration to apply the control is sampled with $T_{max}$ being the maximum duration a control can be applied. The control input is then iteratively applied by numerically integrating the robot dynamics forward in time by a single time step, while simultaneously checking for collisions, until the sampled number of steps has been taken. The condition on line 8 of Algorithm 6 ensures that if the robot is steered into an obstacle, the progress made thus far is not retained. This precaution ensures that no future extensions of the tree originate from a vertex that would immediately encounter an obstacle. If no conflicts are encountered, the newly generated node is returned to be added to the RRT. The kinodynamic RRT continues to expand until either a maximum number of states have been added or the goal region has been reached.
Algorithm 5 Kinodynamic RRT
1:Initial State $x_{0}$
2:$V\leftarrow{x_{0}},E\leftarrow\emptyset,V\leftarrow(V,E)$
3:for $i=1:N$
4:    $x_{rand}\leftarrow\textsc{Sample}()$
5:    $x_{near}\leftarrow\textsc{Nearest}(V,x_{rand})$
6:    $x_{new}\leftarrow\textsc{KinodynamicSteer}(x_{near},x_{rand})$
7:    if $x_{new}\neq NULL$
8:         $V\leftarrow V\cup{x_{new}}$
9:         $E\leftarrow E\cup{(x_{near},x_{new})}$
10:         $G\leftarrow(V,E)$
11:         if $x_{new}\in GoalRegion$  return $G$              return $\emptyset$
Algorithm 6 KinodynamicSteer
1:Closest state in tree $x_{near}$, Target state $x_{rand}$, Robot dynamics function $f$, Maximum control duration $T_{max}$
2:$u\leftarrow\textsc{SampleControl}(x_{near},x_{rand})$
3:$k\leftarrow\textsc{SampleDuration}(T_{max})$
4:$t\leftarrow 0$
5:$x_{new}\leftarrow x_{near}$
6:while $t\leq k$
7:    $x_{new}\leftarrow\textsc{Integrate}(f(x_{new},u))$
8:    $collision\leftarrow\textsc{InCollision}(x_{new})$
9:    if $collision$  return $NULL$
10:    $t\leftarrow t+\Delta t$ return $x_{new}$
Figure 6: In the Warehouse environment, robots located on the top and bottom of each narrow aisle must swap places with the robot they are vertically aligned with.
(a) Tunnels Environment
(b) Tunnels Query
Figure 7: In the Tunnels environment, robots must find and traverse multiple tunnels while moving from the red start positions at the back to the blue goal positions at the front.
(a) GridMaze
(b) GridMaze Query
Figure 8: In the GridMaze environment, robots must move between randomly located starts (red) and goals (blue) within the maze, some of which overlap.
(a) Mining
(b) Mining Query
Figure 9: In the mining environment, robots must swap places with another robot located in an adjacent mine shaft, query locations are shown in red.
@@ V-A Experimental Setup
We measure the performance of Workspace Guided-DaSH against CDR-RRT [10] as well as a set of other approaches that are relevant for comparison. Composite RRT [14] operates in the full composite space, providing a high level of coordination. MRdRRT [17] was originally intended for manipulators, but its implicit search of the composite space makes it pertinent for comparison. SRRT [20] uses subdimensional expansion to limit the size of the search space which is similar in concept to our hypergraph representation. We use Conflict-based Search [36] over the workspace skeleton to perform the guided search for Workspace Guided-DaSH. Each method was implemented in the open-source Parasol Planning Library. All holonomic experiments were run using a computer with an AMD Ryzen 9 7940HS CPU at 4GHz and 32GB of RAM. We run 15 seeds for each scenario. Each method is given 600 seconds to solve a scenario or is considered a failure. We report planning time as well as path cost as given by the makespan.
Figure 10: Average runtime results for the Warehouse scenario. The standard deviation is shown by the shaded region around each line. Data is omitted where all seeds failed.
@@ V-B Environments
V-B1 Warehouse (Fig. 6)
The Warehouse environment features a series of narrow aisles. Robots start on the top and bottom of the aisles and must swap places with the robot with which they are vertically aligned. In the holonomic scenario, the width of the narrow passages is approximately 2.5 times the diameter of the disk robots we consider. We additionally run a kinodynamic variant of this scenario with doubled narrow passage widths and rectangular robots (see Sec. V-D). This scenario measures each approach’s ability to efficiently handle narrow passages within the environment. We highlight Workspace Guided-DaSH’s ability to decompose the problem using topological guidance.
V-B2 Tunnels (Fig. 7)
The Tunnels environment is three-dimensional and features several narrow passages that the robots must traverse. Because of the three-dimensional workspace, the size of the planning space becomes very large. This scenario is intended to measure each algorithm’s performance with respect to the geometric complexity, rather than the congestion, of the environment. We consider L-shaped robots with the width of their longest dimension slightly less than half the width of the narrow passages. Each robot must find and traverse multiple narrow passages to reach its goal, but its shortest path does not conflict with any other robot’s.
Figure 11: Average runtime results for the Tunnels scenario. The standard deviation is shown by the shaded region around each line. Data is omitted where all seeds failed. Data is omitted where all seeds failed.
@@
V-B3 Mining (Fig. 9)
The Mining environment is a 3D environment meant to mimic a set of mine shafts. Again, it features several narrow passages; however, here robots must move in close proximity to each other. We use the same robots as the Tunnels environment. We show that Workspace Guided-DaSH is capable of efficiently finding paths for large robot teams in real-world scenarios.
V-B4 GridMaze (Fig. 8)
@@
Figure 12: Average runtime results for the Mining scenario. The standard deviation is shown by the shaded region around each line. Data is omitted where all seeds failed.
@@
V-C Workspace Guided-DaSH Results
In the Warehouse scenario (results in Fig. 10 and Tab. I), all methods except MRdRRT achieve a 100% success rate in the 2-robot scenario. However, the methods that do not leverage topological guidance begin to struggle with the 4-robot scenario and most cannot solve the 8-robot scenario. CDR-RRT and SRRT are able to solve up to the 8-robot scenario where the exponential size of the composite space leads to a significant decrease in performance. Workspace Guided-DaSH, like CDR-RRT, sees the benefit of using the skeleton representation to guide planning through narrow passages and the coordination provided by using the composite space of multiple robots. Additionally, by using the workspace to inform the planner when to move between different levels of composition, Workspace Guided-DaSH is able to focus planning within smaller spaces than CDR-RRT, making it able to plan for significantly larger robot teams, up to 128 robots, within the 600-second time limit.
In the Tunnels environment (results in Fig. 11 and Tab. II), the size of the search space is increased significantly relative to the Warehouse environment since the robot now has six degrees of freedom. Composite RRT and MRdRRT are unable to find a solution for the 2-robot scenario. This is due both to the large size of the search space and the complexity of the environment. SRRT performs better because of its ability to decompose the search space, but struggles with the narrow passages and only one seed was able to solve the 4-robot scenario. Here, the narrow passage problem is exacerbated since the query requires the planner to find and traverse multiple narrow passages. By leveraging skeleton guidance, CDR-RRT and Workspace Guided-DaSH are able to find solutions for larger groups of robots. Searching the composite space, which grows exponentially with the number of robots, hinders CDR-RRT’s performance, leaving it unable to solve the 64-robot scenario within the time limit. Workspace Guided-DaSH is able to efficiently find solutions for up to 64 robots within the time limit due to its ability to decompose the planning space.
Figure 13: Average runtime results for the GridMaze scenario. The standard deviation is shown by the shaded region around each line. Data is omitted where all seeds failed. The line for Composite RRT is not visible since it only succeeded for the 2-robot scenario.
@@ V-C Workspace Guided-DaSH Results
In the Mining environment (results in Fig. 12 and Tab. III), Composite RRT and MRdRRT are unable to find a solution for the 2-robot scenario due to the large size of the search space and inter-robot conflicts. Similarly, SRRT struggles to find solutions for the 2-robot scenario and cannot solve the 4-robot scenario. By leveraging topological guidance to limit exploration of the composite space, CDR-RRT is able to solve up to the 16-robot scenario. However, it is unable to solve the 32-robot scenario. Workspace Guided-DaSH, by considering the composite space only when necessary, is able to efficiently solve up to the 64-robot scenario within the time limit.
@@
In the 3D GridMaze scenario (results in Fig. 13 and Tab. IV), as a result of the large size of the planning space as well as the difficulty posed by the long curved narrow passages, Composite RRT struggles with the 2-robot scenario and fails to produce a solution for the 4-robot scenario. MRdRRT fails to solve the 2-robot scenario. While CDR-RRT is able to solve the 4-robot scenario, the large size of the composite space prevents it from finding a solution for the 8-robot scenario within the time limit. SRRT benefits from searching a lower-dimensional space when robots are not in conflict, but fails to solve the 32-robot scenario where conflicts become prevalent due to the high level of congestion of the environment. Similar to SRRT, Workspace Guided-DaSH searches lower-dimensional spaces, however, our method also uses the workspace skeleton for conflict resolution. When a conflict is too difficult to be resolved quickly via random sampling, we impose constraints on the robots’ abstract paths through the workspace which allows us to find conflict-free paths more efficiently. Thus, Workspace Guided-DaSH is able to find solutions up to 32 robots. Although this replanning mechanism sometimes results in slightly higher solution costs relative to other methods, a post-processing trajectory optimization step is generally used in practice to smooth paths.
V-D Kinodynamic Experiment Setup
To show the ability of our method to operate on problems with kinodynamic constraints, we use the Kinodynamic RRT 5 as our RRT in the motion hypergraph construction of Workspace Guided-DaSH. Our kinodynamic steering function selects the best control out of $50$ randomly sampled options and has a maximum applied duration of $50$ timesteps.
Figure 14: Average runtime results for the kinodynamic Warehouse scenario. The darker colors indicate seeds that exceeded the time limit of 2700 seconds. Lighter colors show the runtimes of seeds that succeeded to find a path.
@@ V-D Kinodynamic Experiment Setup
We compare the performance of our Kinodynamic Workspace Guided-DaSH against Kinodynamic CBS [29] as it is an unguided multi-robot kinodynamic motion planning method which uses Conflict-Based Search [36] as a high-level planner to resolve conflicts with kinodynamic RRT as the low-level planner. The dynamics are integrated using the $4$th-order Runge-Kutta (RK4) method [37]. K-CBS utilizes a merge bound to inform the high-level search of robots which are repeatedly in conflict with each other after re-planning, drawn from an extension of CBS: Meta-agent CBS [38]. Once this merge bound is reached for a set of robots, the planner merges the robots into a composite robot. The idea is that, if the robots are constantly in conflict with each other, then their coordination is tightly coupled and planning in the composite space will enable the planner to capture the level of coordination needed to find a solution for the set of robots. The authors of Meta-agent CBS [38], state that if the environment consists of narrow corridors where high levels of coordination are required, the merge bound should be set low so the planner can quickly increase the level of coordination being considered to match the problem difficulty. Following his insight, we use a merge bound of $3$ in our experiments to allow the planner to quickly match the level of coordination needed for the challenging problems we are considering. Both methods were implemented in the open-source Parasol Planning Library and experiments were run using a computer with an Intel i9-14900K CPU at 6GHz and 64GB of RAM. Each method was given $45$ minutes to find a solution to each problem.
As in [29], all robots use $2$nd-order car dynamics:
$\dot{x}=v\cos{\theta},\ \dot{y}=v\sin{\theta},\ \dot{\theta}=\frac{v}{l}\tan{\phi},\ \dot{v}=a,\ \dot{\phi}=\gamma$ (2)
with $x$ and $y$ defining the position of the robot and $\theta$ defining the orientation. We let $v$ define the velocity, $\phi$ define the steering angle of the front wheels of the car. $a$ is the first control input representing acceleration, and $\gamma$ as the second control input representing the steering rate. Additionally, we keep the same robot shapes as in [29] but scale down the size of the robots to fit the environment, rectangular robots with length $l=0.35$ and width $w=0.25$.
V-E Kinodynamic Workspace Guided-DaSH Results
As detailed in Sec. V-B1, we run a kinodynamic variant of the Warehouse scenario. In order to account for the increased problem difficulty imposed by kinodynamic constraints, we increased the size of the narrow passages in the environment. Results are shown in Figure 14 and Table V.
In the kinodynamic warehouse environment, K-WG-DaSH can find solutions for problems with up to 6 robots fairly consistently while K-CBS begins to struggle at 4 robots. In the 2 robot scenario, K-WG-DaSH maintains a high success rate with a lower planning time at the expense of a higher path cost as opposed to K-CBS which is inconsistent at finding solutions with a higher planning time but with a lower path cost. At 4 robots, K-CBS sees a significant drop off in success rate, while K-WG-DaSH maintains it’s high success rate. When 6 robots are considered, K-CBS has an extremely difficult time finding solutions and when 8 or 10 robots are considered, it cannot find any solutions in the time limit. K-WG-DaSH sees a slight drop off in success rate at 6 robots, and a larger drop off at 8 and 10 robots, however, it is able to find at least one solution. We argue that the addition of guidance over when to coordinate and topological guidance providing where to go is the reason for the improved success rate and scalability of K-WG-DaSH over K-CBS.
@@ VI Conclusion and Future Work
In this work, we present both the Guided-DaSH framework for hybrid multi-robot planning and an application of this framework to mobile robot MRMP in congested environments. Our adapted framework builds off of a state-of-the-art hybrid planning framework, DaSH, using guidance to enable its application to a wider range of multi-robot planning problems which lack the highly structured search space seen in previous DaSH methods. Additionally, we add support for kinodynamic planning via a modified conflict resolution structure, which also boosts the scalability of our framework by eliminating unnecessary work when finding motion solutions.
Our application of Guided-DaSH to mobile robot MRMP in congested environments, Workspace Guided-DaSH leverages topological guidance in multiple ways, both to inform how to move between different levels of composition during planning and to guide sampling through narrow passages in the workspace. By using knowledge of the workspace, our method is able to efficiently find paths for robot groups of size up to an order of magnitude larger than existing state-of-the-art methods. Building off of this work, we plan to explore the use of different kinodynamic low-level motion planners which plan more intelligently for multi-robot systems by interleaving sampling-based planning and trajectory optimization [39]. Additionally, we plan to explore using parallel computing to further improve the performance of this method, taking advantage of the hypergraph’s ability to naturally decompose the multi-robot planning problems.

### v3

@@ IV-C Kinodynamic Planning
Incorporating kinodynamic constraints into Workspace Guided-DaSH is straightforward due to the modularity of the framework. As kinodynamic constraints influence the motion of the robots, it is imperative that the low-level motion hypergraph, $\mathcal{H}_{m}$, incorporates the kinodynamic constraints during construction. To this end, we adapt the use of CDR-RRT, employed during the construction of $\mathcal{H}_{m}$ to incorporate kinodynamic constraints. This is achieved by substituting their underlying RRT with Kinodynamic RRT [15].
A control input is generated by sampling the control space uniformly at random which is then integrated forward during the RRT extension phase for a randomly sampled duration. In order to utilize the guidance provided by the CDR-RRT, we generate and integrate forward many random control inputs and select the input which moves the system closest to the dynamic sampling region. The kinodynamic RRT continues to expand until either a maximum number of states have been added or the goal region has been reached. We refer the readers to [15] for more details on kinodynamic RRT.
(a) Warehouse
(b) Tunnels Environment
(c) Mining
(d) GridMaze
(e) GridMaze Query
Figure 3: (a) In the Warehouse environment, robots located on the top and bottom of each narrow aisle must swap places with the robot they are vertically aligned with. (b) In the Tunnels environment, robots must find and traverse multiple tunnels while moving from start positions in the back tunnel to goal positions in the front tunnel. (c) In the mining environment, robots must swap places with another robot located in an adjacent mine shaft, robot start positions are located at the front of the tunnels and at intersections between tunnels. (d-e) In the GridMaze environment, robots must move between randomly located starts (red) and goals (blue) within the maze, some of which overlap.
@@ V-A Experimental Setup
We measure the performance of Workspace Guided-DaSH against CDR-RRT [6] as well as a set of other approaches that are relevant for comparison. Composite RRT [14] operates in the full composite space, providing a high level of coordination. MRdRRT [16] was originally intended for manipulators, but its implicit search of the composite space makes it pertinent for comparison. sRRT [19] uses subdimensional expansion to limit the size of the search space which is similar in concept to our hypergraph representation. ARC [20] composes robots into higher dimensional planning spaces based on encountered conflicts, whereas our method uses workspace guidance to preemptively avoid likely conflicts. Thus, comparing our method against ARC effectively measures the impact of this guidance. We use Conflict-based Search [25] over the workspace skeleton to perform the heuristic search for Workspace Guided-DaSH. Each method was implemented in the open-source Parasol Planning Library. Holonomic experiments were run on a 4-core CPU (4GHz, 32GB RAM), while kinodynamic experiments used a 24-core CPU (6GHz, 64GB RAM). We run 15 seeds for each scenario. Each method is given 600 seconds to solve a scenario or is considered a failure. We report planning time as well as path cost given by the makespan.
@@ V-B Environments
V-B1 Warehouse (Fig. 3(a))
The Warehouse environment features several narrow aisles through which robots must swap places with the robot opposite them. In the holonomic scenario, the width of the narrow passages is approximately 2.5 times the diameter of the disk robots we consider. We additionally run a kinodynamic variant with doubled aisle widths. This scenario measures each approach’s ability to efficiently handle narrow passages within the environment. We highlight Workspace Guided-DaSH’s ability to decompose the problem using topological guidance.
V-B2 Tunnels (Fig. 3(b))
The Tunnels environment again features several narrow passages. Here, because of the 3D workspace, the planning space becomes very large. This scenario is intended to measure each algorithm’s performance with respect to the geometric complexity, rather than the congestion, of the environment. We consider L-shaped robots with the width of their longest dimension slightly less than half the width of the narrow passages. Each robot must find and traverse multiple narrow passages to reach its goal, but its shortest path does not conflict with any other robot’s.
@@
V-B3 Mining (Fig. 3(c))
The Mining environment is a 3D environment meant to mimic a set of mine shafts. Again, it features several narrow passages; however, here robots must move in close proximity to each other. We use the same robots as the Tunnels environment. We show that Workspace Guided-DaSH is capable of efficiently finding paths for large robot teams in problems representative of real-world scenarios.
V-B4 GridMaze (Fig. 3(d))
@@
V-C Experimental Results
In the Warehouse scenario (results in Tab. I), all methods except MRdRRT achieve a 100% success rate in the 2-robot scenario. However, the methods that do not leverage topological guidance begin to struggle with the 4-robot scenario and multiple cannot solve the 8-robot scenario. Of those, only ARC is able to solve up to the 32-robot scenario; however, with a success rate of only 26.6%. CDR-RRT and sRRT are able to solve up to the 8-robot scenario where the exponential size of the composite space leads to a significant decrease in performance. Workspace Guided-DaSH, like CDR-RRT, sees the benefit of using the skeleton representation to guide planning through narrow passages and the coordination provided by using the composite space of multiple robots. Additionally, by using the workspace to inform the planner when to move between different levels of composition, Workspace Guided-DaSH is able to focus planning within smaller spaces than CDR-RRT, making it able to plan for significantly larger robot teams, up to 128 robots, within the 600-second time limit.
In the Tunnels environment (results in Tab. II), the size of the search space is increased significantly relative to the Warehouse environment since the robot now has six degrees of freedom. Composite RRT, MRdRRT, and ARC are unable to find a solution for the 2-robot scenario due both to the large size of the search space and the complexity of the environment. sRRT performs better because of its ability to decompose the search space, but struggles with the narrow passages and only one seed was able to solve the 4-robot scenario. Here, the narrow passage problem is exacerbated since the query requires the planner to find and traverse multiple narrow passages. By leveraging skeleton guidance, CDR-RRT and Workspace Guided-DaSH are able to find solutions for larger groups of robots. Searching the composite space hinders CDR-RRT’s performance, leaving it unable to solve the 64-robot scenario within the time limit. Workspace Guided-DaSH is able to efficiently find solutions for up to 64 robots within the time limit due to its ability to decompose the planning space.
@@ V-C Experimental Results
In the Mining environment (results in Tab. III), Composite RRT and MRdRRT are unable to find a solution for the 2-robot scenario due to the large size of the search space and inter-robot conflicts. Similar to the Tunnels scenario, ARC’s performance is impacted by the environment complexity since it does not use topological guidance and it fails to find solutions for the 2-robot scenario within the time limit. sRRT struggles to find solutions for the 2-robot scenario and cannot solve the 4-robot scenario. By leveraging topological guidance to limit exploration of the composite space, CDR-RRT is able to solve up to the 16-robot scenario. However, it is unable to solve the 32-robot scenario. Workspace Guided-DaSH, by considering the composite space only when necessary, is able to efficiently solve up to the 64-robot scenario.
@@
In the 3D GridMaze scenario (results in Tab. IV), as a result of the large size of the planning space as well as the difficulty posed by the long curved narrow passages, Composite RRT struggles with the 2-robot scenario and fails to produce a solution for the 4-robot scenario. MRdRRT fails to solve the 2-robot scenario. While CDR-RRT is able to solve the 4-robot scenario, the large size of the composite space prevents it from finding a solution for the 8-robot scenario within the time limit. sRRT and ARC benefit from searching a lower-dimensional space when robots are not in conflict, but fail to solve the 32-robot scenario where conflicts become prevalent due to the high level of congestion of the environment. Similar to sRRT, Workspace Guided-DaSH searches lower-dimensional spaces, however, our method also uses the workspace skeleton for conflict resolution. When a conflict is too difficult to be resolved quickly via random sampling, we impose constraints on the robots’ abstract paths through the workspace which allows us to find conflict-free paths more efficiently. Thus, Workspace Guided-DaSH is able to find solutions up to 32 robots. Although this replanning mechanism sometimes results in slightly higher solution costs relative to other methods, a post-processing trajectory optimization step is generally used in practice to smooth paths.
V-D Kinodynamic Experiments
To evaluate Workspace Guided-DaSH with kinodynamic constraints, we use Kinodynamic Workspace Guided-DaSH (K-WG-DaSH) incorporating the modifications in Section IV-C. Kinodynamic-CBS [21] (K-CBS) is utilized as our baseline kinodynamic planner as it uses the same Kinodynamic RRT we employ in K-WG-DaSH, however they utilize a merge bound [26] creating composite robots when more coordination is needed. This will allow us to evaluate the benefits of the coordination guidance provided by K-WG-DaSH. Each method was given $45$ minutes to find a solution to planning problems in the warehouse environment, described in Section V-B1, using robots with $2$nd-order car dynamics.
In the kinodynamic warehouse environment (results in Tab.  V), K-WG-DaSH finds solutions for problems with up to 6 robots consistently while K-CBS begins to struggle at 4 robots. In the 2 robot scenario, K-WG-DaSH maintains a high success rate with a lower planning time at the expense of a higher path cost as opposed to K-CBS which is inconsistent at finding solutions with a higher planning time but with a lower path cost. We attribute the scalability to the topological and coordination guidance provided by K-WG-DaSH.
@@ VI Conclusion and Future Work
In this work, we present both the Guided-DaSH framework for hybrid multi-robot planning and Workspace Guided-DaSH, an application of this framework to mobile robot MRMP in congested environments. Our adapted framework builds off of a state-of-the-art hybrid planning framework, DaSH, using guidance to enable its use for a wider range of multi-robot planning problems which lack a highly structured search space. Additionally, we add support for kinodynamic planning via a modified conflict resolution structure, which also boosts the scalability of our framework by eliminating unnecessary work when finding motion solutions.
Workspace Guided-DaSH leverages topological guidance in multiple ways, both to move between different levels of composition during planning and to guide sampling through narrow passages in the workspace. By using knowledge of the workspace, our method is able to efficiently find paths for robot groups of size up to an order of magnitude larger than existing state-of-the-art methods. Building off of this work, we plan to explore the use of different kinodynamic low-level motion planners which plan more intelligently for multi-robot systems by interleaving sampling-based planning and trajectory optimization [27]. Additionally, we plan to explore using parallel computing to further improve the performance of this method, taking advantage of the hypergraph’s ability to naturally decompose the multi-robot planning problems.